OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

Well from this equation we know the slope is -4.. For a line perpindicular to it the slope would be the opposite reciprocal. So it would be 1/4 (basically flip the fraction over the fraction bar and change its sign of negative to posituive or positive to neagtive)... Now plug the new slope 1/4 and point (-16,-11) into the slope intercept formula of: y-y1=m(x-x2) where m is the slope and (x1,Y1) is a point. Does that make sense?

OpenStudy (anonymous):

yes im just not sure how to write out the equation at all.

OpenStudy (anonymous):

y- -11=1/4x- -16?

OpenStudy (anonymous):

very close you have the left side right... Here y- -11= 1/4(x- -16) So this equals y+11=1/4(x+16) because two negatives equal a positive Now notice how the 1/4 is outside the paranthesees. that means we must distribute (or multiply) the 1/4 to everything in the paranthesis. So this equals.. y+11=1/4x+4.. This is youre answer but if you want it in y= something form then just solve for y

OpenStudy (anonymous):

thank you very much!

OpenStudy (anonymous):

Youre very welcome! If you need any more help dont be hesitant to ask!